OpenStudy (welshfella):

Find the next number in the series 0, 0, 3, 20, 115 I've spent about 30 minutes on it but can't get it.

OpenStudy (whpalmer4):

Here's a clue: 1-1 = 0 2-2 = 0 6-3 = 3 24-4 = 20

OpenStudy (welshfella):

AH! Very good! Factorials! 1! - 1 = 0 2! -2 = 0 etc

OpenStudy (welshfella):

what might have made that more difficult is if they had included 0! which is 0! - 0 = 1 1 , 0, 0, 3, 20 would seem very odd!

OpenStudy (whpalmer4):

bingo. there's an online dictionary of integer sequences, btw. https://oeis.org/A005096

OpenStudy (welshfella):

ok thanx

OpenStudy (whpalmer4):

if ever you are having trouble with insomnia, reading a few pages of it ought to do the trick :-)

OpenStudy (welshfella):

lol! yeah!